What and Define The Hamming Metric with example in mathematical foundations of computer science

1
Expert's answer
2021-12-27T16:30:18-0500

For the vectors "a=(a_1,a_2,\ldots,a_n)\in\{0,1\}^n" and "b=(b_1,b_2,\ldots,b_n)\in\{0,1\}^n" the Hamming Metric "\rho_H" is defined as

"\rho_H(a,b)=\sum\limits_{i=1}^n|a_i-b_i|."


For example, for "a=(1,0,1,1)\in\{0,1\}^4" and "b=(0,0,1,0)\in\{0,1\}^4" the Hamming Metric "\rho_H(a,b)" is equal to


"\rho_H(a,b)=|1-0|+|0-0|+|1-1|+|1-0|=1+0+0+1=2."